# Classical JAK2V617F+ Myeloproliferative Neoplasms emergence and development based on real life incidence and mathematical modeling

### v2

@@ 3.3 Age-dependency of JAK2V617F mutation emergence
Let us recall the method. Let $X$ be the vector that corresponds to the observed data of detection ages and has distribution $f_{X}(x;\theta)$ for an age $x$, where $\theta$ is the vector of parameters. Consider the missing data $z$, corresponding to the values of $T_{2}$ with distribution $f_{Z}(z;\theta)$. Then the complete data is $(x,z)$, and it is distributed according to $f_{X,Z}(x,z;\theta)$. If the complete data was available, the the log-likelihood could be calculated by
$\log L_{X,Z}(\theta)=\log f_{X,Z}(x,z;\theta).$
@@
$Q(\theta;\theta_{k})=\mathbb{E}_{\theta_{k}}(\log L_{X,Z}(\theta)|x).$
@@
For the M-step, as suggested in [Rai1993], we take $\theta_{k+1}$ as one Newton-Raphson step from $\theta_{k}$ over the function $Q(\theta_{k+1};\theta_{k})$, that is
@@
and $0<a_{k}\leqslant 1$. The choice of $a_{k}$ needs to be such that it stays in the parameters space and the likelihood is nondecreasing. To do this, take $a_{k}^{0}=1$ and divide by half its value until the parameters satisfy the desired conditions. Then, we apply backtracking line search : while $Q(\theta_{k}+a_{k}\delta_{k};\theta_{k})>Q(\theta_{k};\theta_{k})+10^{-4}a_{k}\nabla Q(\theta_{k};\theta_{k})^{T}\delta_{k}$, set $a_{k}=0.8a_{k}$. Then, we take $a_{k}$ as the final value of these steps.
Once the parameters estimation was done, we performed a Chi-square goodness-of-fit test to study the hypothesis that the data fits both of the two models. We found that the Weibull model was rejected with a p-value smaller than $10^{-17}$ and the Gompertz model was not rejected with a power of $0.85$, indicating that this model seems to properly explain the data. The Gompertz model has been widely used to model different biological age-dependent rates such as mortality [Finch1990] or cancer death rate [Riffenburgh2001]. In our case, it can be explained as follows: we assume that individuals have an initial mutation occurrence rate $A$, and an age-related increase in mutation occurrence represented through a a parameter $k$. Then, for an age $t$, the active mutation occurrence takes the form $\delta_{G}(t)$. Under this hypothesis, the distribution of the time of active mutation emergence $T_{1}$ is the following
$\widehat{f}_{1}(t)=A\,\exp\left(\frac{A}{k}\right)\,\exp\left(k\,t\right)\,\exp\left(-\frac{A}{k}\,\exp(k\,t)\right).$
From (5), we obtain the following distribution for the age of detection $T_{M}$
$\widehat{f}_{M}(t)=\int_{0}^{t}A\,\exp\left(\frac{A}{k}\right)\,\exp\left(k\,s\right)\,\exp\left(-\frac{A}{k}\,\exp(k\,s)\right)\ \frac{1}{(t-s)\sigma\sqrt{2\pi}}\,\exp\left(-\frac{\log(t-s)-\mu}{2\sigma^{2}}\right)ds,$ (6)
@@
since $T_{1}$ and $T_{2}$ are assumed to be independent. By applying the EM-algorithm to our data, we obtain the following estimations on the parameters of the model
$A=0.000122,\quad k=0.095814,\quad\nu=2.083590,\quad\sigma=0.099585.$
With these parameters, we obtain that the mean of $T_{2}$ is $8.07$ years with a standard deviation of $0.81$ years.
3.4 Age-dependency of JAK2V617F mutation emergence with different models of tumor growth
We obtained in Section 3.3 that the Gompertz model with random expansion time is not rejected. But given the small variance of $T_{2}$, we want to investigate wether a fixed expansion time is sufficient to properly explain the data. When considering $T_{2}=\alpha$, we obtain the following distribution for the age of detection $T_{M}$
$\widehat{f}_{M}(t)=A\,\exp\left(\frac{A}{k}\right)\,\exp(k(t-\alpha))\exp\left(-\frac{A}{k}\,\exp(k(t-\alpha)\right).$ (7)
The parameters were estimated using the same EM-algorithm as previously and we obtained the following values
@@ 3.4 Age-dependency of JAK2V617F mutation emergence with different models of tumor growth
Using this estimation, a Chi-square goodness-of-fit test resulted in not rejecting this model with a power of $0.952$. We refer to Figure 5 for visual comparison between the estimations for fixed $T_{2}$ in A-B and random $T_{2}$ in C-D.
Since both models with a Gompertz function for the active mutation occurrence and with a fixed or lognormal expansion time $T_{2}$ are not rejected, we use the Bayesian information criterion (BIC) presented in [Schwarz1978] to select the most appropriate model amongst these two to explain the data. This criterion is defined by
$\text{BIC}=k\log n-2\log\mathcal{L},$
where $k$ is the number of parameters of the model, $n$ is the size of the data sample and $\mathcal{L}$ is the likelihood function. The BIC uses the likelihood function to measure how well a model can explain the data, while also penalizing the number of parameters in the model to avoid overfitting. A lower BIC is preferred. We obtain that the Gompertz model with constant MPN expansion time (7) ($T_{2}$ as a fixed value) has a BIC of 2079 and the Gompertz model with lognormal MPN expansion time (6) has one of 2085.3. This indicates that the model with fixed expansion time should be chosen. While incorporating variability on the expansion time $T_{2}$ gives a better fit to the data in terms of the likelihood, the gain is small and not compensated by the cost of adding additional parameter to the model. This would imply that, given our data, the Gompertz model with constant MPN expansion time is the most appropriate to explain the data. This suggests that environmental modifications (or other sources of variability) do not have a huge impact on the growth phase of the tumor. Using the estimation of the selected Gompertz model with fixed expansion time, we obtained that the expected occurrence time $T_{1}$ is of $63$ years with a standard deviation of $13$ years and the pathology development when the first JAK2V617F stem cell is present, fixed and ready to proliferate is $8.8$ years ($T_{2}$), implying a mean detection age of 76 years.
3.5 Validation of the Gompertz model with fixed expansion time in an external cohort
The French group of MPN recently developed a national database on ET and PV (BCB-FIMBANK: French National Cancer Institute (INCa) BCB 2013 and 2022, CHU Angers) (Table 2). We considered that this database is representative of the general population given the data size ($1111$ JAK2V617F MPN patients) (see Figure 1 C-D for the histogram before and after death adjustment), even if it is not as exhaustive as the previous database that reported all the cases in the region. We used this second data set to validate the Gompertz model with fixed expansion time (7). We performed an estimation of the parameters through the same EM algorithm as previously described and then applied a Chi-Square goodness-of-fit test, which was not rejected with a power of 0.99 (see Figure 6). This proposed model seems to properly explain MPN emergence and proliferation given our two data sets. This suggests that not only the mutation rate cannot be constant through life, but increases exponentially with age following the Gompertz model. Given these positive results to fit the data, we used the Gompertz model with fixed expansion time to compare the differences in mutation acquisition and expansion between the two pathologies PV an ET.
3.6 Differential analysis of Essential thrombocytemia and polycythemia Vera
We hypothesized that ET and PV have a different kinetic because of the double hit in PV compared to single hit in ET: homologous recombination or development of secondary mutations that induces the polycythemia instead of the thrombocytosis only. Given the high number of reported MPN in the FIMBANK cohort, we use this data set to analyze the difference between ET and PV patients.
@@ 3.6 Differential analysis of Essential thrombocytemia and polycythemia Vera
We started by testing if ET and PV corrected data come from the same distribution through a Chi-square test, which rejected this hypothesis with a p-value of $10^{-11}$. Hence, we confirmed that the diagnostic age distributions are different for ET and PV (at a 99% confidence level).
We proceeded to estimate the parameters of each pathology under the Gompertz model with fixed expansion time (7) and found the following values for the parameters, the mean ($m_{1}$) and standard deviation ($s_{1}$) of $T_{1}$ for each disease
@@
The estimation under our model indicates that the average mutation occurrence time ($T_{1}$) is approximately 1.5 years higher for PV which could be explained by the time taken for the mitotic recombination to happen.
@@ 4 Discussion
Our present work provides mathematical modeling of JAK2V617F MPN emergence and expansion based on real life incidence and age at diagnosis, which highlights the role of an exponential age dependence increase of mutation occurrence in MPN. Based on the regional exhaustive registry of Côte d’Or (France), and lastly validated in a MPN national registry (BCB-FIMBANK registry), we modelled the occurrence and expansion times of JAK2V617F MPN. Most of the previous works performed to decipher the time of JAK2V617F “first” cell appearance, are based on genetic hierarchy and back-tracking performed on ancient blood samples or bone marrow samples (see [Osorio2018], [VanEgeren2021], [Mitchell2022]). These genetic approaches rely upon premises of a rate (a frequency) of mutation occurrence as yet unknown in MPN. It necessitates old blood or bone marrow samples for genetic analysis (to recapitulate the genetic tree evolution of the initial cancer cell) which is obviously very rare and probably biased [Hirsch2016]. Genetic analysis inference performed in such approaches are based on mutation appearance but not necessary proliferative/oncogenic. Indeed, it is now well documented that such JAK2V617F mutations could stay for long time without development of pathologies, as seen in clonal hematopoiesis of indeterminate potential (CHIP). Moreover, these genetic back-tracking approaches were built on stable mutation frequencies that overestimate the time of emergence if mutation rate during life is not constant. Our approach was based on registry analysis, differing from classical genetic back-tracking of mutation emergence and we showed that not only the mutation rate cannot be constant, but needs to grow exponentially with age, following the well known Gompertz model. We assumed that two periods of time were necessary to develop and diagnose MPN: a first time period ($T_{1}$) corresponding to the mutation acquisition (genetic event from the conception of the embryo) but also its activation (ie ensuring to develop). This first period of time can correspond biologically to a cell cycle entry of a long-lasting quiescent cell for example, but also includes the possibility of a MPN cancer cell to disappear due to the cancer cell extinction probability. The second time period ($T_{2}$) corresponds to the growth and development of such cancer stem cell until diagnosis. This type of modeling has been applied to invasion-fixation of mutants in eco-evolutionary models (see [Champagnat2006], [Billiard2017]) and to the infectious diseases propagation (see [Barbour2015], [Barbour2013]) but less frequently to cancer emergence and development. Most reported modeling assumes that mutation and invasion rate are aging independent (ie a fixed rate of mutation and activation and a fixed probability of emergence of the pathology for a given patient), which would imply that $T_{1}$ follows an exponential law. However, models under this hypothesis show a peak of cases in younger ages, which is not coherent with the data, where most cases are present for older ages. We then speculated that myeloproliferative neoplasm emergence could need an age-dependent modeling for $T_{1}$. Age dependence does not necessarily mean that the JAK2V617F mutation acquisition is higher in aged stem cells than in young ones; it also could be interpreted as a mutation appearance during childhood or even before birth, but an activation due to external stimulus long time after. The second time (expansion time) is easier to characterize. A model considering the expansion time as a fixed value properly fits the data and is chosen over one as a random variable. This suggests that external changes (stress, food, toxic exposure….) could not significantly modify the behavior of MPN development and that the proliferation time from one “active” cancer cell to the total mass of cancer at diagnosis is relatively fixed. This could mean that the main determining factor is the proliferative speed related to the mutation acquisition. We then can speculate about the possibility (and risks and benefits) to treat subjects with JAK2V617F clonal hematopoiesis to delay the emergence of the MPN or prevent the classical thrombotic complications related to these pathologies. Our mathematical and statistical study allows us to assess the time of emergence and development of MPN. We found that the mean time for cancer cell to emerge is 63 years and time from tumor emergence to diagnosis is approximately 8.8 years in accordance with previous hypothesis, (see [Radivoyevitch2012] and [McKerrell2017]). Based on the Variant Allele Frequency of JAK2V617F in the “FIMBANK” cohort, we observe that the mean JAK2V617F VAF in MPN is around 30% (19.9% in ET and 46.1% in PV respectively) and an increased ratio of 1.4% per year (as reported previously in untreated [Stein2010] or Hydroxyurea only treated patients [Kiladjian2022]. This should mean a 10-year period of time from undetectable VAF to the 30% allele ratio observed at diagnosis in our MPN cohort. This is very close (and a little bit higher) to the 8.8 years of tumor expansion time to MPN diagnosis that we calculated in our model. Furthermore, our modeling fits well with the JAK2V617F CHIP that has been described to take approximatively a decade to transform to MPN [Nielsen2011], [Nielsen2013], [vanZeventer2023]. In conclusion, we demonstrated using mathematical modelling that the emergence and fixation of JAK2V617F mutation is strongly linked to aging mechanisms and that the expansion time of MPN can be considered as non random. Our numerical results show a mean of 63.1 years for the emergence and fixation of the mutation, and an expansion time of approximately 8.8 years. Altogether, these results highlight the interest to test all 50-60 years-old inhabitants for JAK2V617F CHIP and the place of preventive therapies for such subjects.
Tables
Age PMF PV JAK2+ ET JAK2+ Total JAK2+ cases
0 to 4 0 0 0 0
5 to 9 0 0 0 0
10 to 14 0 0 0 0
15 to 19 0 0 0 0
20 to 24 0 0 0 0
25 to 29 0 0 0 0
30 to 34 0 1 4 5
35 to 39 0 0 4 4
40 to 44 0 1 4 5
45 to 49 0 5 4 9
50 to 54 0 2 7 9
55 to 59 1 8 11 20
60 to 64 1 10 6 17
65 to 69 2 13 15 30
70 to 74 4 12 14 30
75 to 79 5 14 29 48
80 to 85 4 15 30 49
More than 85 4 11 23 38
Total number 21 92 151 264
Table 1: First data set (Côte d’Or).
Age PV JAK2+ ET JAK2+ Total JAK2+ cases
0 to 4 0 0 0
5 to 9 0 0 0
10 to 14 0 0 0
15 to 19 2 3 5
20 to 24 1 4 5
25 to 29 5 3 8
30 to 34 8 6 14
35 to 39 18 10 28
40 to 44 20 26 46
45 to 49 48 23 71
50 to 54 57 24 81
55 to 59 78 30 108
60 to 64 80 51 131
65 to 69 114 53 167
70 to 74 113 59 172
75 to 79 84 36 120
80 to 84 69 36 105
85 to 90 30 9 39
90 to 94 7 4 11
Total number 734 377 $11$
Table 2: Second data set (FIMBANK).
Number
of cases
Mean age
at diagnosis
Standard
deviation of age
VAF JAK2
at diagnosis
Standard
deviation of VAF
PV Female 314 66.7 13.7 45.3 22.9
Male 421 64.0 13.3 46.7 30.3
ET Female 229 63.0 15.9 19.2 5.7
Male 147 63.2 13.8 20.9 15.8
Table 3: Information concerning the pathologies PV versus ET by sex.

### v3

@@ 3.2 Parameter estimation
We estimate the parameters using the generalized Expectation-Maximization (EM) algorithm (see for example [12]) as presented in [20]. Let us recall the method. Let $T_{M}$ be the vector that corresponds to the observed data of detection ages and has distribution $f_{T_{M}}(t_{M};\theta)$ for an age $t_{M}$, where $\theta$ is the vector of parameters ($\theta=(A,k,\alpha)$ for Model A.1 (5) and $\theta=(A,k,\mu,\sigma^{2})$ for Model A.2 (6)). Consider the missing data $t_{2}$, corresponding to the values of $T_{2}$ with distribution $f_{T_{2}}(t_{2};\theta)$. Then the complete data is $(t_{M},t_{2})$, and it is distributed according to $f_{T_{M},T_{2}}(t_{M},t_{2};\theta)$. If the complete data was available, the log-likelihood could be calculated by
$\log L_{T_{M},T_{2}}(\theta)=\log f_{T_{M},T_{2}}(t_{m},t_{2};\theta).$
@@
$Q(\theta;\theta_{k})=\mathbb{E}_{\theta_{k}}(\log L_{T_{M},T_{2}}(\theta)|t_{M}).$
@@
For the M-step, as suggested in [20], we take $\theta_{k+1}$ as one Newton-Raphson step from $\theta_{k}$ over the function $Q(\theta_{k+1};\theta_{k})$, that is
@@
and $0<a_{k}\leqslant 1$. The choice of $a_{k}$ needs to be such that it stays in the parameters space and the likelihood is non-decreasing. To do this, take $a_{k}^{0}=1$ and divide by half its value until the parameters satisfy the desired conditions. Then, we apply backtracking line search: while $Q(\theta_{k}+a_{k}\delta_{k};\theta_{k})>Q(\theta_{k};\theta_{k})+10^{-4}\,a_{k}\,\nabla Q(\theta_{k};\theta_{k})^{T}\,\delta_{k}$, set $a_{k}=0.8\,a_{k}$. Lastly, we take $a_{k}$ as the final value of these steps.
3.3 Model selection
We want to compare and select a model amongst Model A.1 (5) and Model A.2 (6). To do so, we estimate the parameters for both models using the EM-algorithm described in subsection 3.2 using the Côte d’Or cohort, and we proceed to perform a Chi-square goodness-of-fit test as follows. We consider the $k=13$ bins corresponding to ages 0300-30, 5 year increments from age $30$ to $85$ and $85+$. For each bin $i$ ($i=1,\dots,k$), let $O_{i}$ and $E_{i}$ be the number and the expected number of observations on bin $i$, respectively. Then, we construct the statistic as
$\chi^{2}=\sum_{i=1}^{k}\frac{(O_{i}-E_{i})^{2}}{E_{i}},$
which follows approximately a Chi-square distribution with kck-c degrees of freedom, where $c$ is the number of estimated parameters ($c=3$ for Model A.1 and $c=4$ for Model A.2).
These tests resulted in not rejecting Model A.1 with a power of $0.952$ and not rejecting Model A.2 with a power of $0.85$. We refer to Figure 2 for visual comparison between the estimations for fixed $T_{2}$ in A-B and random $T_{2}$ in C-D.
Since both models are not rejected, we use the Bayesian information criterion (BIC) presented in [23] to select the most appropriate model amongst these two to explain the data. This criterion is defined by
$\text{BIC}=c\log N-2\log\mathcal{L},$
where $c$ is the number of parameters of the model, $N$ is the size of the data sample and $\mathcal{L}$ is the likelihood function. The BIC uses the likelihood function to measure how well a model can explain the data, while also penalizing the number of parameters in the model to avoid overfitting. A lower BIC is preferred. We obtain that the Model A.1 with constant MPN expansion time ($T_{2}$ as a fixed value) has a BIC of 2079 and Model A.2 with lognormal MPN expansion time has one of 2085.3. This indicates that the Model A.1 with fixed expansion time should be chosen. While incorporating variability on the expansion time $T_{2}$ gives a better fit to the data in terms of the likelihood, the gain is small and not compensated by the cost of adding additional parameter to the model. This would imply that, given our data, the Gompertz model with constant MPN growing time is the most appropriate to explain the data. This suggests that environmental modifications (or other sources of variability) do not have a huge impact on the growth phase of the tumor.
3.4 Model validation
3.4.1 Validation on a second cohort
The French group of MPN recently developed a national database on ET and PV (BCB-FIMBANK: French National Cancer Institute (INCa) BCB 2013 and 2022, CHU Angers) (Table 2). We considered that this database is representative of the general population given the data size ($1111$ JAK2V617F MPN patients) (see Figure 1 C-D for the histogram before and after death adjustment), even if it is not as exhaustive as the previous database that reported all the cases in the region. We used this second data set to validate the Gompertz model with fixed expansion time given by Model A.1 (5). We performed an estimation of the parameters through the same EM algorithm as previously described and then applied a Chi-Square goodness-of-fit test, which was not rejected with a power of $0.99$ (see Figure 3). This model seems to properly explain MPN emergence and proliferation given our two data sets.
3.4.2 Age-independent models and Weibull age-dependent model
The proposed age-dependent model was selected among others given its performance with the data. Indeed, when testing different models for the emergence and expansion time ($T_{1}$ and $T_{2}$), the models fail to fit the data. We study three models with a constant active mutation rate $\delta$, which implies $T_{1}\sim Exp(\delta)$ (see subsection 2.2). This assumption of a constant mutation rate has been taken in the majority of genetic analysis backtracking reported to date [18], [28], [13]. The first and most simple model considers a fixed expansion time $T_{2}$. A second model considers that external factors may influence the expansion time, $T_{2}$ and is hence taken as a lognormal random variable as defined in (2). The third model assumes a fixed expansion time, but allows the time of mutation occurrence to vary from one patient to another or the mutation to develop at different times from one patient to another (quiescence/cycling). This takes into consideration that the biological hypothesis of stress (or environment) influences cancer cell emergence. This model considers $T_{1}$ as an exponential random variable, but its rate is different from one patient to another. Then, for $1\leq i\leq N$, where $N$ is the number of patients, we considered that each age of emergence $T_{1}^{i}$ is drawn from an exponential distribution with parameter $\delta_{i}$. The $\delta_{i}$’s are considered as an independent sample of a $lognormal(m,s^{2})$. The three age-independent models can be summarized as
Model B.1 $:\begin{cases}T_{1}\sim Exp(\delta)\\ T_{2}=\alpha\text{ fixed},\end{cases}$ (7)
Model B.2 $:\begin{cases}T_{1}\sim Exp(\delta)\\ T_{2}\sim lognormal(\mu,\sigma^{2}),\end{cases}$ (8)
Model B.3 $:\begin{cases}T_{1}^{i}\text{ sampled from }Exp(\delta_{i})\\ \delta_{i}\text{ sampled from }lognormal(m,s^{2}),\ i=1,\dots,N\\ T_{2}=\alpha\text{ fixed}.\end{cases}$ (9)
We estimate the parameters of Model B.1 and the ones of Model B.2 through a least-squared estimation which minimizes the sum of the square difference between the data and the model for each age. The parameters of Model B.3 are estimated using the software Monolix [10]. This software is well adapted for this mixed effect modeling as it allows to easily estimate the parameters when considering the survival analysis approach. The estimation is based of the SAEM algorithm and gives robust and global convergence. As illustrated in Figure 4 A, B and E, the approximation for Model B.1 and Model B.3 presents such a big discrepancy with the data that it is sufficient to reject this model. While Model B.2 seems closer to the data, we perform a Chi-square goodness-of-fit test which results in rejecting this model at a $99\%$ confidence level with a p-value smaller than $10^{-17}$. The rejection of all three of these models imply that even when including individual or environmental variations, models with constant active mutation occurrence $\delta$ cannot explain the data.
Figure 4: A: Frequency of cases for the data (in black) and probability density function of the estimation (in red) for the age-independent model (Model B.1 (7)). B: Accumulated frequencies of cases for the data from the (black) and cumulative distribution of the estimation (in red) for the age-independent model (Model B.1 (7)). C: Frequency of cases for the data (in black) and probability density function of the estimation (in red) for the age-independent model with lognormal MPN growing time (Model B.2 (8)). D: Accumulated frequencies of cases for the data (in black) and cumulative density of the estimation (in red) for the age-independent model with lognormal MPN growing time (Model B.2 (8)). E: Accumulated frequencies of cases for the data (in black) and $90\%$ confidence interval of the estimation (in red) for the age-independent model with individual variability (Model B.3 (9))
We can then think of other age-dependent models, such as a Weibull model for the active mutation rate (see [10], pages 90-92). For an age $t$, the Weibull mutation occurrence rate function is given by
$\delta_{W}(t)$ $=\frac{\gamma}{\lambda}\left(\frac{t}{\lambda}\right)^{\gamma-1},$
where $\gamma>0$, $\lambda>0$. We consider the broader model with $T_{2}$ as a lognormal random variable with density function (2). The addition of variability to the expansion time $T_{2}$ should give the same fit or better than a fixed value. The parameters are estimated using the EM-algorithm in subsection 3.2. Then, we perform a Chi-square goodness-of fit test and find that the Weibull model is rejected with a p-value smaller than $10^{-17}$.
Hence, we obtain a good fit of Model A.1 with Gompertz age-dependent active mutation rate and fixed expansion time in both data sets. Coupled with the rejection of all three age-independent mutation rate models and the Weibull age-dependent model, these results suggest that not only the mutation rate cannot be constant through life, but increases exponentially with age following the Gompertz model. Given these positive results to fit the data, we apply Model A.1 (5) to our data to obtain information on the emergence and expansion of the disease and to compare the differences in mutation acquisition and expansion between the two pathologies PV an ET.
4 Applications of the model
4.1 Mutation emergence and expansion times
@@
By estimating the parameters of Model A.1 (5) to the Côte d’Or data set, using the EM-algorithm (subsection 3.2), we obtain the following values
@@
This implies that the expected occurrence time $T_{1}$ is of $63$ years with a standard deviation of $13$ years and the pathology development when the first JAK2V617F stem cell is present, fixed and ready to proliferate is $8.8$ years ($T_{2}$), with a mean detection age of $76$ years.
4.2 Differential analysis of Essential Thrombocythemia and Polycythemia Vera
We hypothesize that ET and PV have a different kinetic because of the double hit in PV compared to single hit in ET: homologous recombination or development of secondary mutations that induces the polycythemia instead of the thrombocytosis only. Given the high number of reported MPN in the FIMBANK cohort, we use this data set to analyze the difference between ET and PV patients.
@@ 4.2 Differential analysis of Essential Thrombocythemia and Polycythemia Vera
We start by testing if ET and PV corrected data come from the same distribution through a Chi-square test, which rejects this hypothesis with a p-value of $10^{-11}$. Hence, we confirm that the diagnostic age distributions are different for ET and PV (at a 99% confidence level).
We proceed to estimate the parameters of each pathology under Model A.1 (5) and find the following values for the parameters, the mean ($m_{1}$) and standard deviation ($s_{1}$) of $T_{1}$ for each disease
@@
The estimation under our model indicates that the average mutation occurrence time ($T_{1}$) is approximately 1.5 years higher for PV which can be explained by the time taken for the mitotic recombination to happen.
@@ 5 Discussion
Our present work provides mathematical modeling of JAK2V617F MPN emergence and expansion based on real life incidence and age at diagnosis, which highlights the role of an exponential age dependence increase of mutation occurrence in MPN. Based on the regional exhaustive registry of Côte d’Or (France), and lastly validated in a MPN national registry (BCB-FIMBANK registry), we model the occurrence and expansion times of JAK2V617F MPN. Most of the previous works performed to decipher the time of JAK2V617F “first” cell appearance, are based on genetic hierarchy and back-tracking performed on ancient blood samples or bone marrow samples (see [18], [28], [13]). These genetic approaches rely upon premises of a rate (a frequency) of mutation occurrence as yet unknown in MPN. It necessitates old blood or bone marrow samples for genetic analysis (to recapitulate the genetic tree evolution of the initial cancer cell) which is obviously very rare and probably biased [7]. Genetic analysis inference performed in such approaches are based on mutation appearance but not necessary proliferative/oncogenic. Indeed, it is now well documented that such JAK2V617F mutations could stay for long time without development of pathologies, as seen in clonal hematopoiesis of indeterminate potential (CHIP). Moreover, these genetic back-tracking approaches were built on stable mutation frequencies that overestimate the time of emergence if mutation rate during life is not constant. Our approach is based on registry analysis, differing from classical genetic back-tracking of mutation emergence and we show that not only the mutation rate cannot be constant, but needs to grow exponentially with age, following the well-known Gompertz model. We assume that two periods of time are necessary to develop and diagnose MPN: a first time period ($T_{1}$) corresponding to the mutation acquisition (genetic event from the conception of the embryo) but also its activation (ie ensuring to develop). This first period of time can correspond biologically to a cell cycle entry of a long-lasting quiescent cell for example, but also includes the possibility of a MPN cancer cell to disappear due to the cancer cell extinction probability. The second time period ($T_{2}$) corresponds to the growth and development of such cancer stem cell until diagnosis. This type of modeling has been applied to invasion-fixation of mutants in eco-evolutionary models (see [4], [3]) and to the infectious diseases propagation (see [1], [2]) but less frequently to cancer emergence and development. Most reported modeling assumes that mutation and invasion rate are aging independent (ie a fixed rate of mutation and activation and a fixed probability of emergence of the pathology for a given patient), which would imply that $T_{1}$ follows an exponential law. However, models under this hypothesis show a peak of cases in younger ages, which is not coherent with the data, where most cases are present for older ages. We then speculate that myeloproliferative neoplasm emergence could need an age-dependent modeling for $T_{1}$. Age dependence does not necessarily mean that the JAK2V617F mutation acquisition is higher in aged stem cells than in young ones; it also could be interpreted as a mutation appearance during childhood or even before birth, but an activation due to external stimulus long time after. The second time (expansion time) is easier to characterize. A model considering the expansion time as a fixed value properly fits the data and is chosen over one as a random variable. This suggests that external changes (stress, food, toxic exposure, …) could not significantly modify the behavior of MPN development and that the proliferation time from one “active” cancer cell to the total mass of cancer at diagnosis is relatively fixed. This could mean that the main determining factor is the proliferative speed related to the mutation acquisition. We then can speculate about the possibility (and risks and benefits) to treat subjects with JAK2V617F clonal hematopoiesis to delay the emergence of the MPN or prevent the classical thrombotic complications related to these pathologies. Our mathematical and statistical study allows us to assess the time of emergence and development of MPN. We find that the mean time for cancer cell to emerge is 63 years and time from tumor emergence to diagnosis is approximately 8.8 years in accordance with previous hypothesis, (see [19] and [11]). Based on the Variant Allele Frequency of JAK2V617F in the “FIMBANK” cohort, we observe that the mean JAK2V617F VAF in MPN is around 30% (19.9% in ET and 46.1% in PV respectively) and an increased ratio of 1.4% per year (as reported previously in untreated [25] or Hydroxyurea only treated patients [9]. This should mean a 10-year period of time from undetectable VAF to the 30% allele ratio observed at diagnosis in our MPN cohort. This is very close (and a little bit higher) to the 8.8 years of tumor expansion time to MPN diagnosis that we calculate in our model. Furthermore, our modeling fits well with the JAK2V617F CHIP that has been described to take approximatively a decade to transform to MPN [16], [17], [29]. In conclusion, we demonstrate using mathematical modelling that the emergence and fixation of JAK2V617F mutation is strongly linked to aging mechanisms and that the expansion time of MPN can be considered as non random. Our numerical results show a mean of 63.1 years for the emergence and fixation of the mutation, and an expansion time of approximately 8.8 years. Altogether, these results highlight the interest to test all 50-60 years-old inhabitants for JAK2V617F CHIP and the place of preventive therapies for such subjects.